Solve:
x^2-y^2=36
x^2-y^2=36

Answer by stanbon(75887)

Solve:
x^2-y^2=36
x^2-y^2=36
--------------------
Solution:
All points on the hyperbola x^2-y^2 = 36
===========================================
Cheers,
Stan H.